OpenStudy (anonymous):

Solve: 0 = 4x2 − 8x − 5. x = −1 over 2, x = 5 over 2 x = −1 over 2, x = −negative 5 over 2 x = 1 over 2, x = −negative 5 over 2 x = 1 over 2, x = 5 over 2

OpenStudy (anonymous):

a

OpenStudy (haichi):

D

Elsa213 (elsa213):

http://openstudy.com/study#/updates/4f9c4405e4b000ae9ed159c9 4x^2-8x-5=0 4x^2+2x-10x-5=0 2x(2x+1)-5(2x+1)=0 (2x+1)(2x-5)=0 2x+1=0 so x=-1/2 2x-5=0 so x = 5/2

OpenStudy (anonymous):

x=-1/2,5/2

Elsa213 (elsa213):

c;

OpenStudy (anonymous):

Thank you guys and thank you Elsa for the explanation <3